# Augmented Lagrangian methods produce cutting-edge magnetic coils for stellarator fusion reactors

### v1

@@ Abstract
Finding feasible coils for stellarator fusion devices is a crux of realizing this concept for future power plants. The coils must reproduce the physics properties of the target plasma while satisfying myriad engineering constraints. Current efforts struggle to navigate the highly nonconvex optimization landscape and result in suboptimal stellarator coils, and/or they end up spending considerable resources scanning the continuous parameter space. In this work, we present an augmented Lagrangian approach to tackle the ill-posed problem of coil optimization. Our approach yields solutions that out-perform previous Pareto-optimal work and display excellent physics and engineering properties. We conclude by illustrating its effectiveness and versatility by generating coils for five stellarators with very different symmetries and magnetic field shaping. In all cases, we find coil solutions that in various ways outperform published coil sets, including built and operating reactors such as Wendelstein 7-X (W7-X) and the Helically Symmetric eXperiment (HSX).
@@ I Introduction
The stellarator, a toroidal magnetic confinement device, stands as a promising candidate for achieving controlled nuclear fusion. Unlike its axisymmetric counterpart, the tokamak, the stellarator relies on complex, three-dimensional magnetic fields to confine the plasma and therefore does not require the use of induced currents that can lead to plasma instabilities [1, 2, 3]. This also allows for steady-state operation, which provides a clear advantage over other technologies to become an energy source. The confining magnetic fields are generated by a system of external coils, the design of which is a critical determinant of the stellarator’s performance and viability. Mathematically, the coils are arbitrary 3D curves in space, and poor optimization of these can easily produce infeasible shapes. Traditionally, stellarator design is subdivided into two stages: stage I, where the plasma boundary shape is optimized, and stage II, where the coil shapes and currents are optimized. This procedure has been used in essentially all previous stellarator designs, including for designing currently operational experiments such as Wendelstein 7-X and HSX [4, 5, 6]. This two-stage process allows for stage-1 optimization to explore suitable physics properties for fusion reactors by solving the magnetohydrodynamic equations of force balance in fixed-boundary [7, 8, 9, 10]. A separate stage-II optimization allows one to focus the optimization on finding feasible coil sets. The optimization of stellarator coils is a magnetostatic ill-posed inverse problem; a multitude of distinct coil configurations can, in principle, generate similar magnetic fields, yet most of these solutions are impractical from an engineering standpoint [11].
@@
Historically, the stellarator community has explored various approaches to tackle the challenge of finding simple yet accurate coils. These have ranged from employing stochastic optimization algorithms to prevent getting stuck in local minima and providing robust configurations, performing large-scale scans capable of exploring vast parameter spaces, investigating diverse coil topologies, such as modular, helical or dipole coils, and even optimizing the plasma and the coils simultaneously [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Despite these varied efforts, finding robust and efficient coil optimization techniques remains a challenge. Currently, various stellarator equilibrium concepts are being investigated also in the industry [25, 26, 27, 28, 29]. Most of these concepts take advantage of the various flavors of omnigeneity, which is a magnetic field property allowing for the confinement of collisionless trapped particles [30, 31, 32, 33]. This property applies to alpha particles that contain a considerable portion of the energy (3.5 MeV) produced from D-T fusion reactions, essential for plasma heating. Among these flavors being pursued are the quasi-axisymmetric (QA), quasi-helical symmetric (QH) and quasi-isodynamic (QI) configurations [9, 34]. Even with this restriction, there is remarkable freedom in the plasma shape and plasma properties [35, 36]. This represents an additional challenge for the development of a robust coil optimization scheme able to reliably deliver stellarator coil sets independently of the complexity of the magnetic configuration.
The engineering requirements for a fusion reactor include limitations on coil curvature, minimum separation distances between coils, considerations for port access for diagnostics, heating systems, and maintenance, and so forth. These requirements must compete with the necessity of high field accuracy to reproduce the desired plasma target. The engineering constraints are immense; to see this, consider just the electromagnetic forces on a reactor-scale stellarator design. Lion et al. [25] report that for the Stellaris reactor-scale design, in order to keep the mechanical integrity of the stellarator, steel plates are placed between the coils. These can have thicknesses “between 0.7 m and 1 m and could be forged, with the heaviest plate in this design weighing 58 tons. The inner rings [around the inner side of the stellarator], about 0.7 m thick and weighing about 200 tons each, can also be forged”. It follows that a stellarator coil design that possesses even mildly reduced forces (e.g. 30%) could result in a significantly more economically viable stellarator. Furthermore, it is estimated that the original Stellaris configuration will have anisotropic deformations in the order of 15 mm, arising from electromagnetic loads. This can potentially result in degraded field quality, asymmetries in the heat loads of island divertors, and so forth. The immense force requirements in this example are not unique to Stellaris, rather, they are emblematic of the inevitable challenges of reactor-scale coil design for stellarators.
@@
I.1 Contributions of this work
This manuscript implements an augmented Lagrangian method for coil optimization, aimed at tackling this multifaceted challenge. We seek to identify best-to-date stellarator coil designs in terms of the plasma and engineering complexity metrics, while avoiding heuristic and manual tuning of the optimization weights. We prove the value of this new method by finding a substantial set of coil solutions that, in various ways, outperform previously published coil sets. The entirety of this work is implemented and can be reproduced in the SIMSOPT optimization suite [37].
@@ II Optimization Method
Early stellarator coil optimization methods considered a single continuous toroidal surface surrounding the plasma, called a winding surface, where sheet currents could flow and a current potential was well-defined. A solution for the current potential would be found, and then coils would be “cut” from them. This approach has been used to find the Wendelstein 7-X coils and can be found in codes such as NESCOIL or REGCOIL [38, 39]. More advanced winding surface methods are in development, especially for generating good initial coils and for performing single-stage optimizations [20]. However, such an approach constrains the coils to sit on a single pre-defined toroidal surface and therefore limits the exploration of better coil sets. Recent efforts have introduced the concept of filamentary coils that can be described as a truncated Fourier series, allowing for the adjustment of the Fourier coefficients as the optimization is performed, and avoiding any surface restriction. This is the representation used in this paper within the SIMSOPT software framework, as described below in Cartesian coordinates for the i-th coil,
@@
where are the various inequality and equality constraints combined and for are the weights. Inequality constraints are typically recast to equality constraints by , since this quantity vanishes if is below the target value. As was pointed out in Conlin et al. for plasma optimization [40], the weights have to be set by the user, rarely represent an optimal tradeoff, and therefore have to be adjusted. It is also only possible to satisfy the constraints when . Moreover, in order to circumvent the struggle of weight adjustment, large-scale scans of targets/weights are used, which can be computationally expensive and suffer restriction to local regions of the parameter space. In order to tackle these issues, the following augmented Lagrangian formulism for coil optimization is proposed with the Lagrangian,
@@
The term is the standard Lagrange multiplier term, and is the penalty term defining the Hadamard product between and . The latter penalizes constraint violations. Note that what previously were user-determined hyperparameters become now Lagrange multipliers that are meant to be updated throughout the optimization, so time consuming manual updates are avoided. Moreover the penalty term prevents the necessity of infinite magnitudes for the , avoiding the two main disadvantages of the previous formalisms. The algorithm employed to steer and update the Lagrange multipliers is inspired from [40, 41, 42] and it is here presented and adapted for clarity:
@@
Here is the tolerance of the Lagrangian gradient and is the tolerance for the measure of constraint violation, and together they set the convergence condition. The argmin optimization is an internal minimization step where for fixed Lagrange multipliers the local optimizer L-BFGS-B [43] is used, a quasi-Newton method suitable for problems with a high number of degrees of freedom. In the context of SIMSOPT the individual gradients of each constraint function are calculated using a combination of analytic derivatives and automatic differentiation, where the chain rule is applied using vector-Jacobian products allowing for efficient derivative estimations. As mentioned previously, the goal is to design feasible coils that are able to reproduce exactly the magnetic field of fusion plasmas. Therefore, the main objective of the optimization is to minimize the field error. This is achieved via the normalized squared flux, which is a measure of deviation between the coil magnetic field and the target magnetic surface,
@@
Regarding the remaining coil constraints, we use standard metrics for complexity in the literature [44, 45]:
Objective | Bound | Motivation |
Squared Flux | Sufficient accuracy to preserve important plasma properties. | |
Minimum CS distance | m | Sufficient distance for neutron-absorbing blanket. |
Minimum CC distance | m | Sufficient distance for avoiding coil intersections and assembly difficulties from finite coil build. |
Total coil length | m | As small as possible to decrease the costs of purchasing superconducting material. But there must be sufficient coil length for coils to be far enough from the plasma to reduce coil-surface distance and improve squared flux. |
Max curvature | m-1 | Low enough local curvature to simplify geometry and manufacturing costs. |
Max MSC | m-1 | Similar to the curvature , the average curvature must remain low so that the coils can be manufactured precisely without extensive cost. |
Coil linking # | 0 | Interlinking coils could greatly increase the complexity of assembly. |
Max force | MN/m | Max forces must stay within the force tolerances of the HTS material, estimated for industrial REBCO-based cables [46, 47, 48]. |
@@
is a function penalizing the minimum distance between the last closed flux surface of the plasma and the coils, preventing it from being smaller than . This can allow enough space for the positioning of a tritium-breeding blanket and shielding wall to keep the integrity of the coils and fusion device that can be damaged by high energy neutrons. restrains the minimum coil-to-coil distance from being smaller than , accounting for finite dimensions of the coils as well as coil assembly and diagnostics. The length of the coils is set by a quadratic penalty that penalizes the sum of length of all the coils if it goes above the threshold . Longer coils means more expensive superconducting material is required to construct them. The cost function calculates the norm of the curvature of a coil and then penalizes the local curvature of the coil if it exceeds the threshold . Higher curvature coils tend to be more difficult to manufacture and assemble. In order to prevent configurations with interlinked coils, which would be very challenging to build in practice, the Gauss linking number is calculated for each pair and becomes an integer if they are interlinked. Finally, minimizes the pointwise self and mutual forces that are above some threshold representing the tolerances of the superconducting material in the coils. The details for the force estimation can be found in Hurwitz et al. [49]. It has become an increasing necessity to find configurations minimizing the coil-to-coil forces because the required coil currents in a reactor-scale stellarator can reach tens of Mega-Ampere values [50, 51, 20, 52, 45]. For the purpose of this work, the estimation of the maximum force per HTS turn is done by dividing the total force value by the number of HTS turns per coil.
@@
In order to determine the relative HTS material requirements for coil sets with varying number of coils per half-field period, for all the following results in this work we assume 200 HTS turns per coil for a reference coilset (e.g. 800 turns total in a four-coil solution) and recalculate the number of turns assuming fixed total current per half-field period (e.g. a three coil solution will be assumed to have HTS turns in each coil to keep the total current fixed). This is probably a pessimistic lower bound on the amount of HTS required for solutions with fewer coils, but it also helps to mitigate the forces in these designs.
@@ III.1 Quasi-axisymmetric stellarator
Lastly, at a given fixed total length, we attempted some three-coil and five-coil solutions. Five-coil solutions were found with comparable or marginally improved performance to the four-coil solutions, so we concluded that the extra coil per half-field period was not worth the inevitable reduction in plasma access. However, assuming that fairly long modular coils can be constructed in practice, an excellent solution was found with m total length coil solution with only three coils per hfp. It replicates most of the equivalent (i.e. same total coil length) four-coil solution complexity metrics while exhibiting better field accuracy and higher coil-coil distances. It can be seen visually in Fig. 2 to provide very good plasma access. These benefits will have to be weighed against the additional coil turns in the three-coil solution that will be necessary to continue avoiding the critical current limits in the superconducting material.
For completeness, the Boozer plot which displays qualitatively the departure from quasisymmetry together with the Poincaré plots showing the nested flux surface structure of the magnetic equilibrium are shown in the Appendix A. Straight lines tell that very low levels of quasisymmetric error are achieved.
@@ III.2 Quasi-helical symmetric Stellarator
In a second optimization demonstration, we target the plasma boundary of the Landreman&Paul quasi-helically symmetric (QH) configuration [9]. Like quasi-axisymmetric plasmas, QH equilibria are designed to confined collisionless trapped particles having also reduce alpha-particle losses. The QH configuration is also the only type of quasisymmetry that has been effectively built with the Helically Symmetric eXperiment (HSX)[6], which is explored in Section III.5. One of the reasons QH symmetry and more generally quasisymmetric stellarators are of interest for a potential fusion reactor lies in the prediction of shear flows in the plasma, allowing for the suppression of microturbulence [54]. However, a four-field-period QH configuration is considered to be more demanding from a coil optimization standpoint since the plasma contains more shaping than the QA configuration shown previously. The objective is to reproduce particle confinement properties of this configuration to high enough accuracy. Therefore, the coils here obtained are compared and optimized using the same set-up as in [55], where the main purpose was to obtain a suitable coil configuration at reactor-scale. The major radius of this configurations is approximately 13 m, the aspect ratio is approximately 8 and the average rotational transform is 1.24.
@@
In the original work in Wiedman et al. [55], five coils per hfp were chosen as it was considered that fewer coils did not yield high enough accuracy to reproduce the plasma equilibrium at the coil-to-coil and coil-to-surface distances required. Here, we show that not only it is possible to reproduce a similar configuration as in [55], but it is also possible to reduce the number of coils to four per half field period. The total coil length is also reduced by slightly relaxing other physical requirements. The results for both these optimizations are displayed on Table 3. Together the average field accuracy and maximum field error are comparable to the original work for both the four coil and five coil configurations, the stellarators are shown in Figure 3. Except for m-reduced minimum coil-coil distance and requiring more HTS material, the four-coil solution is as good, or better, in every engineering metric as the Wiedman solution.
@@
To further verify that these coils generate magnetic fields that that fulfill physics requirements, we plot the Boozer plot of the outmost closed magnetic surface together with the Poincaré cross-sections of the equilibrium, shown in Appendix A. As expected the magnetic field amplitude contours appear straight with the low-field regions displaying higher levels of coil ripple than as reported by Wiedman et al., however this is consistent with the fact that the configuration has one less coil per half-field period than previously. It is also possible to reveal the nested structure of the QH equilibrium without any apparent magnetic island chains, capturing the high degree of shaping of the equilibrium.
Property | #1 | #2 | Wiedman |
of coils per hfp | 4 | 5 | 5 |
Max | |||
Max Force [MN/m] | 0.36 | 0.31 | |
Total Length [m] | 160 | 177.8 | 177.8 |
Min CS Distance [m] | 1.91 | 1.59 | 1.62 |
Min CC Distance [m] | 0.8 | 1.09 | 1.09 |
Max [] | 1 | 0.77 | 0.81 |
Max MSC [] | 0.1 | 0.08 | 0.079 |
Quasisym. Error | |||
HTS length [km] | 284.5 | 284.5 |
@@
Confinement of collisionless trapped particles is also strongly suggested by the low quasisymmetric error present in these configurations. Note that this is achieved without necessarily breaching the thresholds and activating the constraints in the optimization. In particular, this is visible in table 3 for the coil-surface distance (lower bound set to m) and the maximum curvature (upper bound set to m-1) for the QH4 and QH5 configurations respectively. This shows that it is possible to set loose upper and lower bounds and still obtain feasible solutions that do not necessarily achieve the bounds. In Section III.5, we present results in the same spirit as for this configuration but for the fully manufactured and assembled QH HSX stellarator.

### v2

@@ Abstract
Finding feasible coils for stellarator fusion devices is a critical challenge of realizing this concept for future power plants. Current design efforts struggle to navigate the highly nonconvex optimization landscape, spend considerable resources scanning the parameter space, and may produce suboptimal coils. In this work, we present an augmented Lagrangian approach to tackle the ill-posed problem of coil optimization. We illustrate its effectiveness and versatility by generating coils for five stellarators with very different symmetries and magnetic field shaping. In all cases, we find Pareto-optimal coil solutions that in various ways outperform published coil sets.
@@ I Introduction
The stellarator, a toroidal magnetic confinement device, stands as a promising candidate for achieving controlled nuclear fusion. Unlike its axisymmetric counterpart, the tokamak, the stellarator relies on complex, three-dimensional magnetic fields to confine the plasma and therefore does not require the use of induced currents that can lead to plasma instabilities [1, 2, 3]. This also allows for steady-state operation, which provides a clear advantage over other technologies to become an energy source. The confining magnetic fields are generated by a system of external coils, the design of which is a critical determinant of the stellarator’s performance and viability.
Early stellarators were built from relatively straightforward coils (circular or helical shapes); unfortunately, the plasma confinement turned out to be much worse than anticipated. In recent decades, there has been a resurgence in stellarator research – first in the public and now also in the private sector – that has been driven by advances in theory and optimization, especially with respect to designing the magnetic field itself; this has come at the cost of considerably more complicated coils. Mathematically, these coils are arbitrary 3D curves in space, and current approaches to calculating them can easily produce infeasible shapes.
Traditionally, stellarator design is subdivided into two stages: stage I, where the plasma boundary shape is optimized, and stage II, where the coil shapes and currents are optimized. This procedure has been used in essentially all previous fully developed stellarator designs, including for designing currently operational experiments such as Wendelstein 7-X and HSX [4, 5, 6]. This two-stage process allows for stage-1 optimization to explore suitable physics properties for fusion reactors by solving the magnetohydrodynamic equations of force balance in fixed-boundary [7, 8, 9, 10]. A separate stage-II optimization focuses on finding a corresponding coil set. However, the optimization of stellarator coils is a magnetostatic ill-posed inverse problem; a multitude of distinct coil configurations can, in principle, generate similar magnetic fields, yet most of these solutions are impractical from an engineering standpoint [11].
@@
Historically, the stellarator community has explored various approaches to tackle the challenge of finding simple yet accurate coils. These have ranged from employing stochastic optimization algorithms to prevent getting stuck in local minima and to produce robust configurations, performing large-scale scans capable of exploring vast parameter spaces, investigating diverse coil topologies, such as modular, helical or dipole coils, and even optimizing the plasma and the coils simultaneously [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Despite these varied efforts, finding robust and efficient coil optimization techniques remains a challenge.
Currently, various stellarator equilibrium concepts are being investigated also in the industry [25, 26, 27, 28, 29]. Most of these concepts take advantage of the various flavors of omnigeneity, which is a magnetic field property allowing for the confinement of collisionless trapped particles [30, 31, 32, 33]. This property applies to alpha particles that contain a considerable portion of the energy (3.5 MeV) produced from D-T fusion reactions, essential for plasma heating. Among these flavors being pursued are the quasi-axisymmetric (QA), quasi-helical symmetric (QH) and quasi-isodynamic (QI) configurations [9, 34]. Even with this restriction, there is remarkable freedom in the plasma shape and plasma properties [35, 36]. This represents an additional challenge for the development of a robust coil optimization scheme able to reliably deliver stellarator coil sets independently of the complexity of the magnetic configuration.
Furthermore, it is estimated that the original Stellaris configuration will have anisotropic deformations in the order of 15 mm, arising from electromagnetic loads. This can potentially result in degraded field quality, asymmetries in the heat loads of island divertors, and so forth. The immense force requirements in this example are not unique to Stellaris, rather, they are emblematic of the inevitable challenges of reactor-scale coil design for stellarators.
@@
This manuscript describes the implementation of an augmented Lagrangian method for coil optimization, aimed at tackling this multifaceted challenge. We seek to identify best-to-date stellarator coil designs in terms of the plasma and engineering complexity metrics, while avoiding heuristic and manual tuning of the optimization weights. We demonstrate the value of this new method by finding a substantial set of coil solutions that, in various ways, outperform previously published coil sets. The entirety of this work is implemented and can be reproduced in the SIMSOPT optimization suite [37].
@@ II Optimization Method
II.1 Prior developments in coil optimization
Early stellarator coil optimization methods considered a single continuous toroidal surface surrounding the plasma, called a winding surface, where sheet currents could flow and a current potential was well-defined. A solution for the current potential would be found, and then coils would be “cut” from them. This approach has been used to find the Wendelstein 7-X coils and can be found in codes such as NESCOIL or REGCOIL [38, 39]. More advanced winding surface methods are in development, especially for generating good initial coils and for performing single-stage optimizations [20]. However, such an approach constrains the coils to sit on a single pre-defined toroidal surface and therefore limits the exploration of better coil sets.
Recent efforts have introduced the concept of filamentary coils that can be described as a truncated Fourier series, allowing for the adjustment of the Fourier coefficients as the optimization is performed, and avoiding any surface restriction. This is the representation used in this paper within the SIMSOPT software framework, as described below in Cartesian coordinates for the i-th coil,
@@ II.1 Prior developments in coil optimization
where are the various inequality and equality constraints combined and for are the weights. Inequality constraints are typically recast to equality constraints by , since this quantity vanishes if is below the target value. As was pointed out in Conlin et al. for plasma optimization [40], the weights have to be set by the user, rarely represent an optimal tradeoff, and therefore have to be adjusted. It is also only possible to satisfy the constraints when . Moreover, in order to circumvent the struggle of weight adjustment, large-scale scans of targets/weights are used, which can be computationally expensive and suffer restriction to local regions of the parameter space.
II.2 Augmented Lagrangian formalism
In order to tackle these issues, the following augmented Lagrangian formulism for coil optimization is proposed with the Lagrangian,
@@ II.2 Augmented Lagrangian formalism
The term is the standard Lagrange multiplier term, and is the penalty term defining the Hadamard product between and . The latter penalizes constraint violations. Note that what previously were user-determined hyperparameters become now Lagrange multipliers that automatically updated throughout the optimization, so time consuming manual updates are avoided. Moreover the penalty term prevents the necessity of infinite magnitudes for the , avoiding the two main disadvantages of the previous formalisms. The algorithm employed to steer and update the Lagrange multipliers is inspired from [40, 41, 42] and it is here presented and adapted for clarity:
@@
Here is the tolerance of the Lagrangian gradient and is the tolerance for the measure of constraint violation, and together they set the convergence condition. The argmin optimization is an internal minimization step where for fixed Lagrange multipliers the local optimizer L-BFGS-B [43] is used, a quasi-Newton method suitable for problems with a high number of degrees of freedom. In the context of SIMSOPT the individual gradients of each constraint function are calculated using a combination of analytic derivatives and automatic differentiation, where the chain rule is applied using vector-Jacobian products allowing for efficient derivative estimations.
II.3 Treatment of squared flux
As mentioned previously, the goal is to design feasible coils that are able to reproduce exactly the magnetic field of fusion plasmas. Therefore, the main objective of the optimization is to minimize the field error. This is achieved via the normalized squared flux, which is a measure of deviation between the coil magnetic field and the target magnetic surface,
@@ II.4 Engineering-informed coil constraints
Regarding the definitions of the usual constraints () on the optimization, we use standard metrics for coil complexity, as given in the literature [44, 45]:
| Objective | Bound | Motivation | |||
| Normalized |
| Minimum CS distance | m |
| Minimum CC distance | m |
| Total coil length | m |
| Max curvature | m-1 |
| Max MSC | m-1 |
| Coil linking | 0 |
| Max force | MN/m |
@@
is a function penalizing the minimum distance between the last closed flux surface of the plasma and the coils, preventing it from being smaller than . This can allow enough space for the positioning of a tritium-breeding blanket and shielding wall to keep the integrity of the coils and fusion device that can be damaged by high energy neutrons. restrains the minimum coil-to-coil distance from being smaller than , accounting for finite dimensions of the coils as well as coil assembly and diagnostics. The length of the coils is set by a quadratic penalty that penalizes the sum of length of all the coils if it goes above the threshold . Longer coils means more expensive superconducting material is required to construct them. The cost function calculates the norm of the curvature of a coil and then penalizes the local curvature of the coil if it exceeds the threshold . Similarly, penalizes the mean squared curvature of a curve when it exceeds the threshold , this yields smoother curves. Higher curvature coils tend to be more difficult to manufacture and assemble. In order to prevent configurations with interlinked coils, which would be very challenging to build in practice, the Gauss linking number is calculated for each pair and becomes an integer if they are interlinked. Finally, minimizes the pointwise self and mutual forces that are above some threshold representing the tolerances of the superconducting material in the coils. The details for the force estimation can be found in Hurwitz et al. [49]. It has become an increasing necessity to find configurations minimizing the coil-to-coil forces because the required coil currents in a reactor-scale stellarator can reach tens of Mega-Ampere values [50, 51, 20, 52, 45]. For the purpose of this work, the estimation of the maximum force per HTS turn is done by dividing the total force value by the number of HTS turns per coil.
@@
In order to determine the relative HTS material requirements for coil sets with varying number of coils per half-field period, for all the following results in this work we assume 200 HTS turns per coil for a reference coilset (e.g. 800 turns total in a four-coil solution) and recalculate the number of turns assuming fixed total current per half-field period (e.g. a three coil solution will be assumed to have HTS turns in each coil to keep the total current fixed). This is probably a pessimistic lower bound on the amount of HTS required for solutions with fewer coils, but it also helps to mitigate the forces in these designs. These calculations are also valid if using Low Temperature Superconductors (LTS).
@@
II.5 Comparison with standard methods
In this section, we compare current filamentary methods implemented in SIMSOPT with the augmented Lagrangian approach. For this we use a common continuation method where the coils are optimized with a small number of degrees of freedom, and that solution is used to initialize subsequent optimizations with more and more degrees of freedom. We start with order coils and solve successive optimization problems until . One run therefore corresponds to 16 successive optimizations. In this example we use the Landreman&Paul quasi-axisymmetric plasma equilibrium scaled at 1 meter major radius found in [9] and include all the constraints as specified in Equation 7. For the standard method, as the weights are traditionally user-determined, we choose to perform 20 runs with randomly assigned weights in a small range centered around known close-to-optimal weights from previous work. We then compare these 20 runs to an identical set-up but optimized using the augmented Lagrangian formalism as described in Section II.2. The results are displayed in Figure 1.
The figure illustrates all of the resulting optimization metrics that are penalized in the constraints. These plots show that standard methods are generally unable to balance all the constraints, while the augmented Lagrangian method is able to find a solution within all the upper and lower bounds. Out of the 20 runs, the runs 3, 12, and 13 violate the least number of constraints (only one) but significantly fail this constraint (a prescribed force tolerance). In contrast, the augmented Lagrangian run successfully passes all tests, without needing any weight tuning. We emphasize the strength of this result; without significant manual weight tuning, the standard methods cannot regularly produce solutions with a prescribed list of constraints like the augmented Lagrangian results.
@@ III.1 Quasi-axisymmetric stellarator
Lastly, at a given fixed total length, we attempted some three-coil and five-coil solutions. Five-coil solutions were found with comparable or marginally improved performance to the four-coil solutions, so we concluded that the extra coil per half-field period was not worth the inevitable reduction in plasma access. However, assuming that fairly long modular coils can be constructed in practice, an excellent solution was found with m total length coil solution with only three coils per hfp. It replicates most of the equivalent (i.e. same total coil length) four-coil solution complexity metrics while exhibiting better field accuracy and higher coil-coil distances. It can be seen visually in Fig. 3 to provide very good plasma access. More remarkably, although it was expected that a low number of coils per hfp would reduce the degree of quasisymmetry because of increased coil ripple, the 3-coil solution still manages to outperform the Wechsung L18 solution. Throughout this work, we use the two-term formulation defined in the Appendix A for a quantitative measure of the quasisymmetric error displayed in Table 2. Moreover, a decrease of about 3 orders of magnitude in the quasisymmetric error is visible as the total coil length threshold is increased. These benefits will have to be weighed against the additional coil turns in the three-coil solution that will be necessary to continue avoiding the critical current limits in the superconducting material.
For completeness, the Boozer plot which displays qualitatively the departure from quasisymmetry together with the Poincaré plots showing the nested flux surface structure of the magnetic equilibrium are shown in the Appendix B. Straight lines tell that very low levels of quasisymmetric error are achieved.
@@ III.2 Quasi-helical symmetric Stellarator
In a second optimization demonstration, we target the plasma boundary of the Landreman&Paul quasi-helically symmetric (QH) configuration [9]. Like quasi-axisymmetric plasmas, QH equilibria are designed to confined collisionless trapped particles having also reduce alpha-particle losses. The QH configuration has only been effectively built once with the Helically Symmetric eXperiment (HSX)[6], which is explored in Section III.5. One of the reasons QH symmetry and more generally quasisymmetric stellarators are of interest for a potential fusion reactor lies in the prediction of shear flows in the plasma, allowing for the suppression of microturbulence [54]. However, a four-field-period QH configuration is considered to be more demanding from a coil optimization standpoint since the plasma contains more shaping than the QA configuration shown previously. The objective is to reproduce particle confinement properties of this configuration to high enough accuracy. Therefore, the coils here obtained are compared and optimized using the same set-up as in [55], where the main purpose was to obtain a suitable coil configuration at reactor-scale. The major radius of this configurations is approximately 13 m, the aspect ratio is approximately 8 and the average rotational transform is 1.24.
@@
In the original work in Wiedman et al. [55], five coils per hfp were chosen as it was considered that fewer coils did not yield high enough accuracy to reproduce the plasma equilibrium at the coil-to-coil and coil-to-surface distances required. Here, we show that not only it is possible to reproduce a similar configuration as in [55], but it is also possible to both improve the engineering metrics and reduce the number of coils to four per half field period. The total coil length is also reduced by slightly relaxing other physical requirements. The results for all three optimizations are displayed on Table 3. Together the average field accuracy and maximum field error are comparable to the original work for both the #1 and #2 configurations. The stellarators are shown in Figure 4. Except for m-reduced minimum coil-coil distance and more HTS material, the four-coil solution is as good, or better, in every engineering metric as the Wiedman solution.
@@
To further verify that these coils generate magnetic fields that that fulfill physics requirements, we plot the Boozer plot of the outmost closed magnetic surface together with the Poincaré cross-sections of the equilibrium, shown in Appendix B. As expected the magnetic field amplitude contours appear straight with the low-field regions displaying higher levels of coil ripple than as reported by Wiedman et al., however this is consistent with the fact that the configuration has one less coil per half-field period than previously. It is also possible to reveal the nested structure of the QH equilibrium without any apparent magnetic island chains, capturing the high degree of shaping of the equilibrium.
| Property | #1 | #2 | #3 | Wiedman |
| of coils per hfp | 4 | 5 | 5 | 5 |
| 122 | ||||
| Max | 41.2 | |||
| Max Force [MN/m] | 0.36 | 0.33 | 0.31 | |
| Total Length [m] | 160 | 177.8 | 160 | 177.8 |
| Min CS Distance [m] | 1.91 | 1.59 | 2.0 | 1.62 |
| Min CC Distance [m] | 0.8 | 1.09 | 1.20 | 1.09 |
| Max [] | 1 | 0.77 | 0.39 | 0.81 |
| Max MSC [] | 0.1 | 0.08 | 0.05 | 0.079 |
| Quasisym. Error | 1602 | |||
| HTS length [km] | 284.5 | 256.0 | 284.5 |
@@
Confinement of collisionless trapped particles is also strongly suggested by the low quasisymmetric error present in configurations #1 and #2. Note that this is achieved without necessarily breaching the thresholds and activating the constraints in the optimization. In particular, this is visible in table 3 for the coil-surface distance (lower bound set to m) and the maximum curvature (upper bound set to m-1) for the #1 and #2 configurations respectively. This shows that it is possible to set loose upper and lower bounds and still obtain feasible solutions that do not necessarily achieve the bounds.
Finally, the goal of configuration #3 is to show that by relaxing the field accuracy target our approach is able to improve every engineering metric that is optimized. As shown in table 3, while the field accuracy is in the order of 1%, configuration #3 achieves the all the best engineering properties, except for the forces, which in this case were not directly targeted during optimization. Notably there is a 38cm increase of coil-to-surface distance, a 17.8m decrease in the coil length per hfp, and a 11cm increase in the coil-to-coil distance. This translates into a 10% cost reduction in HTS tape, while having considerably simpler coils with 52% percent less local curvature and 37% percent less mean squared curvature. The simplicity of the coils is visible in Figure 4. The maximum force also remains comparable to the Wiedman case. However, the quasisymmetric error reaches values that are three orders of magnitude higher than the other configurations. Although the field accuracy is still within reasonable bounds, the rapid deterioration of quasisymmetry comes from high sensitivity of the equilibrium to field errors from the coils. As was mentioned previously, the Landreman&Paul configuration was targeted to reach minimal levels of quasisymmetric errors. Complementary, single-stage methods can be employed to improve the compatibility between equilibrium and coils. In Section III.5, we present results in the same spirit as for this configuration but for the fully manufactured and assembled QH HSX stellarator.